### v4

# Efficient first-principles approach to Gibbs free energy with thermal expansion

Abstract
We propose a method to evaluate the Gibbs free energy from constant-volume first-principles phonon calculations. The volume integral of the pressure is performed by determining the volume and the bulk modulus in equilibrium at finite temperatures, where the pressure and its volume derivative are evaluated utilizing first-principles calculations of the Grüneisen parameter without varying the volume. We validate our method for fcc Al by comparing with the conventional quasiharmonic approximation. Furthermore, we integrate our method with self-consistent phonon theory and apply it to calculations for bcc Ti, hcp Ti, and tetragonal ZrO2. We demonstrate the accuracy and computational efficiency of our method by comparing results with those obtained from directly volume-varied self-consistent phonon calculations. In all cases, our method accurately evaluates the free energy change due to thermal expansion using only constant-volume phonon calculations.
I Introduction
The Gibbs free energy is a critical quantity for understanding the thermodynamic properties of materials. Moreover, it serves as the basis for not only phase stability but also nonequilibrium dynamics, such as the evolution of microstructures [1, 2]. Due to its importance, countless experiments have been conducted and compiled into databases over the years, in line with the development of the CALPHAD (CALculation of PHAse Diagrams) method. However, the evaluation of the free energy based solely on experiments has become increasingly difficult with the growing complexity of multicomponent materials. Therefore, there has been a growing interest in thermodynamics assessment through first-principles calculations, particularly using density functional theory (DFT) [3], to address this challenge [4]. Although DFT is valid only at the ground state, finite-temperature properties can be estimated by combining statistical mechanical models with DFT. Additionally, data assimilation, which integrates numerical and experimental data, allows for the high-precision prediction of materials properties, even when experimental data is limited [5, 6]. These high-precision predictions have the potential to accelerate materials development by enhancing the accuracy of free energy calculations.
Various methods exist for incorporating thermodynamic contributions into first-principles free energy assessments: cluster-expansion and variation methods for mixing enthalpy and entropy [7, 8, 9, 10, 11, 12, 13], phonon calculations for vibrational free energy [14, 15], and spin models for magnetic free energy. Furthermore, interactions between different contributions are crucial for the quantitative evaluation of the free energy [16, 17, 18]. For an accurate assessment of the free energy over a wide range of temperatures, the contribution of thermal expansion is also important.
Thermal expansion originates primarily from lattice anharmonicity, at least in nonmagnetic materials. A prominent computational method for evaluating thermal expansion is the quasiharmonic approximation (QHA) [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Within the QHA, the equilibrium volume at a finite temperature is determined by performing harmonic phonon calculations on structural models with varying volumes. However, the QHA is not applicable to high-temperature phases that exhibit lattice instability at low temperatures, due to imaginary frequencies. To restore lattice stability at finite temperatures, it is necessary to consider the anharmonic effects of phonons. Several approaches exist to incorporate anharmonicity, such as temperature-dependent effective potentials [32, 33, 34], piecewise polynomial potential partitioning [35, 36], self-consistent ab initio lattice dynamics [37], and self-consistent phonon (SCPh) theory [38, 39]. Unfortunately, in any of these approaches, the direct modeling of thermal expansion, as in the QHA, is computationally expensive because these methods require substantial resources even for a single-volume calculation. This poses a crucial challenge in the systematic evaluation of free energies for materials with diverse compositions. Previous studies have highlighted the significance of thermal expansion, emphasizing its impact on phase stability and phase-transition temperatures [40, 41, 42, 43, 44, 45, 46, 30]. Therefore, an efficient approach for evaluating the effects of thermal expansion is in demand.
In this study, we present an alternative approach that evaluates the change in the free energy due to thermal expansion from phonon calculations at a single volume. This approach eliminates the need for computationally demanding calculations involving many different volumes of supercells. Our approach is based on the combination of the vibrational pressure and bulk modulus via Grüneisen theory [47]. From phonon calculations at a single volume, we obtain the equilibrium volume and equilibrium bulk modulus, which serve as parameters in the second-order Birch-Murnaghan equation of state at any temperature. Using these parameters, the change in the free energy due to thermal expansion is calculated by integrating pressure with respect to volume. We perform first-principles calculations of the Gibbs free energy and the coefficient of linear thermal expansion for fcc Al, which is a system with large thermal expansion coefficient, to validate our approach. Furthermore, to demonstrate the accuracy and computational efficiency of our approach, we integrate it with SCPh theory and perform calculations of bcc and hcp Ti, as well as tetragonal ZrO2, and compare the results with those obtained from directly volume-varied SCPh calculations.
II Theory
The Gibbs free energy , where is the temperature and is the external pressure, is obtained via a Legendre transformation as
where is the Helmholtz free energy, expressed as a function of and the volume [44], and is the equilibrium volume at and . The mathematical notation of means finding the minimum value of the function by changing the variable . Therefore, can be obtained as
where means the value of variable at which the function is minimized. According to Eq. (1), evaluating typically requires Helmholtz free energies over a range of volumes, since the minimization is performed with respect to volume. Instead, we introduce a practical scheme that reconstructs using only the Helmholtz free energy at a fixed reference volume , i.e., , aided by an equation of state.
II.1 Volume integral of pressure (VIP) method
Our objective is to determine the Gibbs free energy of a single phase, where is a set of numbers of atoms with the element index . In the following discussion, we consider closed systems. Therefore, we do not write as a variable explicitly. The Gibbs free energy can be written as
where is the pressure as a function of and .
To calculate the integral term in Eq. (3), we propose a method named as the volume integral of pressure (VIP) method. Here, we use the second-order Birch-Murnaghan equation of state (EOS) [48, 49, 50]
where is the bulk modulus at . The parameters and can be determined by evaluating and its first derivative
at . The left-hand sides of Eqs. (4) and (5) are numerically calculated from first principles. Thus, and are determined by solving the set of equations. For nonmagnetic materials, the Helmholtz free energy within the Born-Oppenheimer approximation can be divided into the following contributions [51]
where is the internal energy at zero temperature except for the zero-point vibrational energy, is the electronic free energy defined as a change from the zero-temperature case, and is the vibrational free energy. For a given atomic configuration, is calculated as the ground-state total energy from first principles. The electronic contributions to the free energy can be calculated within the finite-temperature DFT [52, 53]. From the relation
the pressure and its volume derivative for each term in Eq. (6) are obtained. Since computational costs of are in general modest, a set of this quantity for different volumes can be used to calculate the contributions to the pressure and its volume derivative . In contrast, the key point of the VIP method is that we evaluate the vibrational contributions to Eqs. (4) and (5) from first-principles phonon calculations at single volume , as explained in the following subsection.
II.2 Vibrational pressure and its volume derivative
We derive analytical formulae for and its volume derivative, to evaluate the left hand sides in Eqs. (4) and (5). First, we express the vibrational pressure in terms of the vibrational free energy as follows:
where is the volume-dependent frequency of the th phonon with the wave vector , is the Boltzmann constant, is the Bose-Einstein distribution function with zero chemical potential, and is the mode Grüneisen parameter [45, 47, 14, 54] as follows:
where is the polarization vector and is the variation of the dynamical matrix due to a volume change. Appendix A contains the details of the definitions. The derivative of with respect to the volume can be written as
II.3 Volume dependence of parameters
If the contribution of the term containing in Eq. (II.2) is minor, we can evaluate the Gibbs free energy without phonon calculations at various volumes. To test this assumption, we redefine Eq. (II.2) as
where and are defined as follows:
Note that Eq. (15) is directly related to the bulk modulus as
To evaluate the impact of the on the Gibbs free energy, we performed calculations for fcc Al, diamond Si, bcc Ti, hcp Ti, and tetragonal ZrO2. These calculations explicitly varied the volume to assess the influence of the on the Gibbs free energy. Our results show that the contribution of the is negligible in all cases. These numerical results are presented in Sec. S2 of the Supplemental Material [55]. Therefore, Eq. (II.2) can be approximated as
From the above discussions, the flowchart of the VIP method is as illustrated in Fig 1. The true strength of the VIP method lies in its seamless integration with SCPh calculations. During the fitting process to obtain fourth-order force constants for SCPh calculations, the third-order force constants are determined simultaneously, thereby eliminating any additional computational costs [38].
II.4 Computational details
We performed first-principles calculations based on density functional theory within the projector augmented wave (PAW) method [56] as implemented in the Vienna simulation package (VASP) [57, 58]. The generalized gradient approximation (GGA) parameterized by Perdew, Burke, and Ernzerhof (PBE) [59] was employed for the exchange-correlation functional. Supercells of 333 (108 atoms) for fcc Al, and 444 (128 atoms) for both hcp and bcc Ti were employed. The cutoff energy was set to 313 eV and 232 eV for Al and Ti, respectively. A -centered -point grid was used for Al, while -centered and grids were used for bcc and hcp Ti, respectively. For tetragonal ZrO2, we used a supercell with 108 atoms and a Monkhorst-Pack -point grid of . The cutoff energy was set to 520 eV and the exchange-correlation functional was treated within the PBEsol [60] according to Refs. [61, 62, 63]. Interatomic force constants (IFCs) and phonon frequencies were calculated using the ALAMODE package [64]. The harmonic IFCs were calculated using the so-called direct method with atomic displacements of 0.01 Å. To calculate mode Grüneisen parameters, third-order IFCs were estimated using the least absolute shrinkage and selection operator (LASSO) technique from randomly displaced structures, as implemented in the ALAMODE package [64]. For hcp Ti, we explicitly optimized the ratio at each temperature. Our calculations reveal that the ratio changes very small, with a maximum variation of only about 0.3 % over the studied temperature range. This small change is consistent with the experimental measurement [65] and the previous computational study [66] and it supports our approximation that thermal expansion along the and axes is nearly identical. The ratio is presented in Sec. S1 of the Supplemental Material [55]. For bcc and hcp Ti, we employed molecular dynamics (AIMD) at 1155 K to generate atomic configurations at finite temperature. Moreover, harmonic IFCs of hcp Ti were extracted from atomic configurations at finite temperature generated by AIMD at 300 K. In the estimations of IFCs, we considered anharmonic terms up to the sixth and fourth order for bcc and hcp Ti, respectively. In addition, for tetragonal ZrO2, we employed AIMD to generate random atomic configurations at 1000 K and performed SCPh calculations including fourth order IFCs. We calculated the Born effective charges and the dielectric tensor of tetragonal ZrO2 using density functional perturbation theory and obtained values of , , , , , , , which agree well with the previous computational result [67]. The vibrational free energy based on SCPh theory is represented in Eq. (8) with an additional correction term [64, 68]. However, this correction term was disregarded when calculating the vibrational pressure and its volume derivative because its contribution is small compared with the harmonic term. Therefore, the vibrational frequency was simply replaced with . The electronic free energy calculated by using the density of state (DOS) and the Fermi-Dirac distribution function as follows [69]:
where
and is the Fermi level. The in Eq. (20) is calculated from the ground-state total energy first-principles calculations. The electronic pressure and its volume derivative were calculated using the second-order Birch-Murnaghan EOS for convenience. Note that the order of the Birch-Murnaghan EOS has almost no effect on the determination of and .
In the calculations for bcc and hcp Ti, to approximately account for electron-phonon coupling effects, we also consider the case of replacing the electronic free energy with that for the atomic configurations with vibrational displacements as follows [69]:
where
and is the temperature-dependent DOS modulated by the vibrational atomic displacements. In this case, we generate vibrational atomic displacement at finite temperature [14] using the ALAMODE package [64] as follows:
where is the number of atoms in the primitive cell, is the mass of the atoms, is the polarization vector, and the normal mode amplitude are randomly sampled from Gaussian distribution [70, 71] as
where is the variance as follows:
We confirmed that quantitative differences between Helmholtz free energies evaluated by Eqs. (6) and (25) are negligible for fcc Al due to strong itineracy of valence electrons. Thus, we used Eq. (6) for fcc Al, whereas Eq. (25) were used for bcc and hcp Ti at . For tetragonal ZrO2, Eq. (6) was employed since the electronic contribution is negligible due to its insulating nature.
III Applications
We apply the VIP method to fcc Al, which is a system with a large coefficient of linear thermal expansion (CLTE), as a validation of our method. To further illustrate the effectiveness and computational efficiency of the VIP method, we combine it with SCPh theory to perform calculations on bcc and hcp Ti, as well as tetragonal ZrO2, and then compare these results with those obtained from directly volume-varied SCPh calculations. We also determine the phase transition point of Ti by Gibbs free energy calculated using the VIP method. In the following calculation we considered zero external pressure (), and in Eq. (3) is chosen as the volume obtained from structural optimization at zero temperature. This selection serves as the reasonable and conventional starting reference point for practical calculations. The sensitivities of the VIP method to the selection of reference volume for fcc Al and Ti are detailed in Sec. S4 of the Supplemental Material [55].
III.1 fcc Al
Figure 2 shows the comparison between the VIP method and the QHA in terms of the free energy and the CLTE for fcc Al over the temperature range from 0 to 1000 K. The term in Eq. (17), which arises from the volume dependence of the Grüneisen parameters, was evaluated by directly varying the volume as a part of the validation of the VIP method. We first evaluate the change in the Helmholtz free energy due to thermal expansion as shown in Fig. 2(a). The change in the Helmholtz free energy is about meV/atom at 1000 K. This indicates that the contribution of thermal expansion to the free energy is significant at high temperatures. We also compared the Gibbs free energy calculated by the VIP method with that obtained by the QHA, as shown in Fig. 2(b). The negligible difference between and suggests that we can accurately predict the change in the Helmholtz free energy without incorporating into the calculation. Our method can calculate the Gibbs free energy within meV/atom of the QHA up to 1000 K, even for fcc Al, where the contribution of thermal expansion is significant. We also calculated the CLTE for fcc Al, as shown in Fig. 2(c). The difference in the CLTE between calculations with and without considering is negligible. The VIP results show a good agreement at lower temperatures, up to 600 K, with a deviation of less than 2%.
However, a discrepancy exists between the results of the VIP method and the experimental data [65] at high temperatures around 1000 K. This discrepancy is attributed to the accuracy of EOS; results obtained using the third-order Birch-Murnaghan EOS within the QHA shown as the green line in Fig. 2(c) significantly reduce the deviation from the experimental data. Although the error arises from the order of the EOS, we emphasize that even in fcc Al — which has a large coefficient of thermal expansion and is therefore highly sensitive to differences in EOS accuracy — this level of error still allows for an adequate description of the coefficient of thermal expansion. We therefore expect that the VIP method will accurately evaluate the CLTE of many other systems with smaller thermal expansion than fcc Al. The details regarding the accuracy of the second-order Birch-Murnaghan EOS for fcc Al are provided in Sec. S3 of the Supplemental Material [55]. Moreover, similar discussions for the VIP method applied to diamond Si are available in Sec. S4 in the Supplemental Material [55].
Differences between the Gibbs energies by DFT and the CALPHAD data are presented in Fig. 3. To examine the temperature dependence, we shifted the origin of the DFT calculations to match the CALPHAD data at 300 K. Figure 3 clearly shows that the difference between the Helmholtz free energy and the CALPHAD data grows rapidly with increasing temperature, while the Gibbs free energy exhibits much better agreement due to inclusion of thermal expansion. This result demonstrates that considering the change in free energy due to thermal expansion is crucial for accurately evaluating phase stability.
III.2 bcc and hcp Ti
Similar to the discussion on fcc Al, the Gibbs free energy can be accurately calculated without including (Sec. S2 in the Supplemental Material [55]). Therefore, in the following calculations, we used to determine the Gibbs free energy with the VIP method. Figure 4 shows the difference in the Gibbs free energy of Ti between bcc and hcp from 300 to 2000 K. The origin of the vertical axis is shifted that the values from first principles and those of CALPHAD coincides at 300 K, because the GGA overestimates the energy of the bcc phase using the static bcc lattice as a local maximum of the potential energy surface [66]. Both the Gibbs free energy and the Helmholtz free energy follow a similar curve until the phase transition temperature. Thermal expansion significantly affects the absolute values of the Gibbs free energy as shown in Sec. S4 of the Supplemental Material [55]. However, because the thermal expansion contributions to the Gibbs free energy in the bcc and hcp phases are very similar, their effects nearly cancel out when calculating the free energy difference. In contrast, considering vibrational atomic displacements significantly affects the evaluation of the relative phase stability; the Gibbs free energy leads to a better agreement with the CALPHAD data by including electron-phonon couplings . Table 1 lists the phase transition temperatures of Ti in the different definitions of free energy.
| Computational conditions | ||
| w/ | 1167 K | |
| w/ | 1171 K | |
| w/ | 1189 K | |
| w/ | 952 K | |
| w/ | 941 K | |
| w/ | 961 K | |
| CALPHAD | 1155 K |
Figure 5 shows the difference in the free energy between the VIP method and the CALPHAD data [72] of bcc Ti from 900 to 2000 K. The DFT values are shifted to match the CALPHAD data at 1150 K that is indicated by the vertical dashed line. As observed with Al, incorporating the change in the free energy due to thermal expansion improves the agreement with the CALPHAD data for including . Even though the Helmholtz free energy , which excludes the effect of appears to reproduce the CALPHAD data better, this agreement can be attributed to merely error cancellation, as thermal expansion and electron-phonon couplings have opposite tendencies with each other as is clear from the figure. Namely, electron-phonon couplings increase the Gibbs free energy of the bcc phase. Particularly, the VIP method accurately reproduces the Gibbs free energies obtained by volume-varied SCPh calculations.
The difference in the electronic free energy between bcc and hcp Ti is shown in Fig. 6. As is seen from the figure, the effect of electron-phonon couplings destabilizes the bcc phase, where the difference in the electronic free energy between bcc and hcp Ti decreases from approximately 50 meV/atom to 10 meV/atom. It should be noted that the maximum errors in the electronic free energy for five vibrational atomic displacements were 0.5 meV/atom for bcc and 0.2 meV/atom for hcp, indicating that even a single sampling may be sufficient.
In connection with the above discussion, Figs. 7(a) and (b) show the electronic DOS of bcc and hcp Ti, respectively. In bcc, the DOS at the Fermi level decreases with increasing temperature. In contrast, the DOS at the Fermi level increases with increasing temperature in hcp. A particularly notable point is that the randomization of atomic coordinates leads to a smoother DOS with increasing temperature. Although the 3 valence electrons of Ti are localized, the randomization of atomic coordinates due to vibrational atomic displacement leads to a smoother DOS and significant changes in the DOS at the Fermi level. The electronic free energy difference between the bcc and hcp phases, caused by changes in the DOS due to vibrational atomic displacements, i.e., electron-phonon couplings, plays a pivotal role in determining the phase transition temperature. Omitting this contribution leads to substantial discrepancies, particularly at elevated temperatures. Finally, we present the computational costs for calculating the Gibbs free energy of bcc and hcp Ti within SCPh theory using the VIP method, as detailed in Table 2, which indicates a significant advantage of our method.
| CPU time (core hour) of DFT calculations | ||
| VIP | volume-varied SCPh | |
| bcc Ti | 243 | 1,953 |
| hcp Ti | 348 | 1,860 |
| tetragonal ZrO2 | 186 | 1,432 |
III.3 tetragonal ZrO2
Figure 8 shows the comparison between the VIP method and the volume-varied SCPh calculations for tetragonal ZrO2 in terms of the free energy and the coefficient of volumetric thermal expansion (CVTE) over the temperature range from 0 to 1500 K. The difference in the Gibbs free energy between the VIP method and the volume-varied SCPh calculations is less than 1 meV/atom over the temperature range from 0 to 1500 K. This figure demonstrates the applicability of the VIP method to accurately and efficiently reproduce the results obtained by volume-varied SCPh calculations even for tetragonal ZrO2 with six atoms in the primitive cell. Calculated CVTE exhibits the deviation of approximately 7.5% from the volume-varied SCPh at 1500 K, but improved accuracy is also observed at lower temperatures (e.g., deviation approximately 4.5% at 600 K). We employed -varied SCPh calculations that independently varied the and axes to minimize the Gibbs free energy at each temperature. The resulting ratio can be considered constant in the temperature range of 0 to 1500 K, with a maximum change of less than 0.06% from 0 K. The results of the -varied SCPh calculations for tetragonal ZrO2 in terms of the CLTE and the CVTE over the temperature range from 0 to 1500 K are shown in Sec. S5 of the Supplemental Material [55]. Table 2 also lists the computational costs for calculating the Gibbs free energy of tetragonal ZrO2 within SCPh theory using the VIP method.
IV Summary
We have proposed a method, i.e., the VIP method, for calculating the Gibbs free energy from constant-volume phonon calculations. A notable advantage of our method compared with the conventional QHA is its ability to calculate vibrational pressure and its volume derivative from single-volume first-principles phonon calculations using Grüneisen parameters resulting in significant reduction in computational costs. We applied our method to fcc Al exhibiting significant thermal expansion, where the Gibbs free energy and the linear coefficient of thermal expansion calculated by the VIP method agree well with the QHA and experimental results. Furthermore, we also applied our method to bcc and hcp Ti, where effects of thermal expansion and electron-phonon couplings nearly cancel out in evaluating the Gibbs free energy of the bcc phase. Considering both effects, the hcp-bcc phase-transition temperature agreed satisfactorily with experimental results. Moreover, we applied the VIP method to tetragonal ZrO2. In these cases, the Gibbs free energy and the coefficient of linear thermal expansion calculated by the VIP method agree well with the volume-varied SCPh calculations. Due to low computational costs of our method, we expect high-throughput first-principles phase diagram calculations will become possible opening up new pathway for materials design considering phase equilibria. Such an approach should also be reinforced by, e.g., data assimilation techniques.
Acknowledgements.
This work was partly supported by JSPS-KAKENHI Grant No. JP24K01144 and by MEXT-DXMag Grant No. JPMXP1122715503. The calculations were partly carried out by using supercomputers at ISSP, The University of Tokyo, and TSUBAME4.0, Institute of Science Tokyo.Appendix A FORCE CONSTAMTS AND GRÜNEISEN PARAMETER
If the vibrational displacements are small compared with the interatomic distance, the potential energy where is the atomic position of the interacting atomic system can be expanded in a power series of the displacements as follows:
where and is the displacement of atom in the th unit cell. In. Eq. (33) the linear term is omitted because atomic forces are zero in equilibrium positions. The coefficient defined as
are called the th order interatomic force constants in real-space representation. The phonon frequency can be obtained by diagonalizing dynamical matrix as
where is the polarization vector and is the atomic mass of atom and is the translational vector of the -th unit cell. Grüneisen parameter in Eq. (12) is derived from the change in the dynamical matrix [54, 74] as
Supplemental Material
Efficient first-principles approach to Gibbs free energy with thermal expansion
Supplemental Materials
Kota Hashimoto
Tomonori Tanaka
Yoshihiro Gohda
Appendix S1 Temperature dependence of the in hcp Ti
Figure S1 illustrates the temperature dependence of the ratio in hcp Ti. The ratio was calculated using the QHA. The maximum change in the ratio from 0 K is about 0.3 %; therefore, approximating the ratio as a constant is reasonable.
Appendix S2 Contributions of for Gibbs free energy
Figure S2 shows the differences in the Gibbs free energy between and , calculated using the VIP method, for fcc Al, diamond Si, bcc Ti, hcp Ti, and tetragonal ZrO2, respectively. These results show that the contribution of is negligible in all cases.
Appendix S3 Convergence of the second-order Birch-Murnaghan equations of state for fcc Al
Figure S3 shows the difference between the fitting curves obtained using the second- and third-order Birch-Murnaghan equations of state for fcc Al. The second-order Birch-Murnaghan EOS converges well onto DFT data around 15 % of the reference volume, which is sufficient for our calculations.
Appendix S4 The sensitivity of the VIP method to the choice of reference volume
We have confirmed the sensitivity to the selection of the reference volume for both fcc Al, which exhibits significant thermal expansion, and bcc Ti, which exhibits pronounced anharmonicity. We used an expanded reference volume of 1.06 (i.e., a 6 % expansion relative to ), where was obtained from structural optimization at zero temperature and zero external pressure. The difference in the Gibbs free energy between the VIP method using and the QHA for fcc Al is shown in Fig. S4. The Gibbs free energy calculated using the VIP method when selecting 1.06 as the reference volume reproduces the QHA results to within 1 meV/atom across the temperature range. Figure S5 (a) shows the difference in the Gibbs free energy between bcc and hcp Ti from 300 to 2000 K. As shown in Fig. S5 (b) and (c), the use of 1.06 changes the Gibbs free energy for the bcc and hcp phases by approximately 2 meV/atom and 1 meV/atom, respectively, over the temperature range. Moreover, this selection has a negligible effect on the evaluation of the phase transition temperature.
Appendix S5 Applications
In this section, we present the results obtained by the VIP method for various materials. We explicitly calculated the term in Eq. (17) by performing calculations at different volumes for benchmarking purposes in the VIP method. These figures show that the VIP method can accurately reproduce the results obtained by the QHA and the volume-varied SCPh calculations.
S5.1 diamond Si
The comparison between the VIP method and the QHA for diamond Si in terms of the free energy and the coefficient of linear thermal expansion (CLTE) are shown in Fig. S6.
S5.2 Ti
S5.2.1 bcc Ti
The comparisons between the VIP method and the volume-varied SCPh calculations for bcc Ti in terms of the free energy and the CLTE are shown in Figs. S7 and S8. To incorporate thermal expansion, we performed SCPh calculations at multiple volumes. Figure S7 considers vibrational atomic displacements in the electronic free energy; Fig. S8 does not.
S5.2.2 hcp Ti
The comparisons between the VIP method and the volume-varied SCPh calculations for hcp Ti in terms of the free energy and the CLTE are shown in Figs. S9 and S10. To incorporate thermal expansion, we performed multiple SCPh calculations at different volumes. Figure S9 considers vibrational atomic displacements in the electronic free energy; Fig. S10 does not.
S5.3 tetragonal ZrO2
Figure S11 shows the coefficient of thermal expansions for tetragonal ZrO2 obtained by the -varied SCPh calculations, which independently varied the and axes to minimize the Gibbs free energy at each temperature. The resulting ratio of 1.443, exhibiting good agreement with the experimental result at 973 K (a deviation of only about 0.02 %) presented in [76]. Furthermore, the ratio can be treated as constant in the temperature range of 0 to 1500 K, given that its maximum change from 0 K is less than 0.06 %. The coefficient of volumetric thermal expansion (CVTE) calculated from the volume-varied SCPh, agrees about 5 % of the result obtained by the -varied SCPh at 1200 K as shown in Fig. S11 (a). In addition, the Gibbs free energies calculated from the volume-varied SCPh reproduce those of -varied SCPh calculations, with the deviation between them reaching a maximum of 0.4 meV/atom at 1500 K and lessening at lower temperatures. Therefore, the volume-varied SCPh calculations can be utilized to accurately determine the Gibbs free energy for tetragonal ZrO2, even without explicitly varying the ratio. For reference, Fig. S11 (b) also shows the CLTEs for the and axes.